# Explicit Visual Prompt Instruction Tuning

### v1

@@ 3.2 Inst-IT Bench
Existing multimodal benchmarks primarily focus on global understanding, failing to provide more in-depth insights into the instance-level comprehension capability of models. To fill this gap, we present Inst-IT Bench, a benchmark specifically designed to diagnose instance-level understanding in both images and videos.
@@
Construction Process. Inst-IT Bench is built through a semi-automated workflow. Initially, we apply the pipeline described in Sec. 3.1 to generate around 20 open-ended QA pairs for each image and video. These generated pairs are then manually reviewed to ensure both accuracy and diversity. Specifically, we remove overly simple questions, ensuring the remaining ones are focused on specific instances. We also refine the questions and answers, making necessary rephrasing to ensure correctness. After this rigorous process, each sample retains 2 to 5 carefully selected question-answer pairs from the original 20, forming the final test set. In addition, we generate three hard negative answers for each question, forming a multiple-choice QA with four options. In total, Inst-IT Bench includes 1,000 QA pairs for 338 images and 1,000 QA pairs for 206 videos. Each QA pair is available in both open-ended and multiple-choice formats. More details can be found in Appendix B.
@@
Metrics. Each test sample in Inst-IT Bench includes both open-ended and multiple-choice formats. For open-ended QAs, we leverage GPT-4o to score the response from a model based on its similarity to the ground-truth answer. For multiple-choice QAs, we calculate the average accuracy across all samples. More details on the metric calculations can be found in Appendix B.
3.3 Inst-IT Dataset
Instruction fine-tuning is crucial for training multimodal models, however, we find there lacks instruction fine-tuning datasets that focus on specific instances. This gap hinders the development of instance-level understanding of the community. Using the data annotation pipeline presented in Sec. 3.1, we create a large-scale instruction-tuning dataset Inst-IT Dataset. To the best of our knowledge, this is the first instruction fine-tuning dataset that provides fine-grained annotations centric on specific instances
Data Sources. We utilize five video instance segmentation datasets (BRUST [3, 17], UVO [75], OVIS [59], LVVIS [71] and YoutubeVIS-2021 [82]) and two object tracking datasets (BenSMOT [35], VidOR [67]) as our video sources, as they provide annotations of instance locations, which is useful in SoM visual prompting [81]. For the image source, we select the SA-1B [27] dataset due to its diversity and abundance of instance objects. Altogether, we collect 51k images and 21k video samples as our visual data pool. More details can be found in Appendix C.
Statistics. Inst-IT Dataset contains 21k videos and 51k images. On average, each video includes one video-level description, 7.3 frame-level descriptions (each covering individual instances, the entire scene, and temporal differences), and 15.6 open-ended question-answer pairs. We treat images as static, single-frame videos without temporal descriptions. In total, Inst-IT Dataset includes visual data from 21k videos and 51k images, along with 21k video-level descriptions, 207k frame-level descriptions, and 335k open-ended QA pairs. For more details about Inst-IT Dataset, please refer to Appendix C.
3.4 Instruction Tuning with Inst-IT Dataset
@@
We adopt the widely-used LLaVA-Next [41] architecture to evaluate the effectiveness of our Inst-IT Dataset. We train our model under an image-video joint training pipeline, where we mix our Inst-IT Dataset with the open-source LLaVA-Next-DATA [43]. For single-image samples, we follow the original AnyRes paradigm [41] to split and encode sub-images according to the aspect ratio. For video and multi-image data, we batch the samples together, encode them, and flatten them into a sequence. Additionally, we apply $2\times 2$ spatial pooling to reduce the number of visual tokens in the video inputs. More details are in Sec. 4.1.
Model LLM #IT Image Video Open-Ended Q&A Multi-Choice Q&A Open-Ended Q&A Multi-Choice Q&A Random Guess - N/A - 25.0 - 25.0 GPT-4o [55] - N/A 74.1 84.8 65.5 81.0 Gemini-1.5-flash [64] - N/A 65.3 79.5 57.9 75.8 Open-source image models LLaVA-1.5 [40] Vicuna-7B [15] 665K 41.6 32.1 - - ViP-LLaVA [5] Vicuna-7B [15] $\sim$1.2M 42.1 29.2 - - SoM-LLaVA [79] Vicuna-7B [15] 695K 45.1 40.0 - - LLaVA-Next [41] Vicuna-7B [15] 765K 46.0 42.4 - - Open-source video models LLaVA-NeXT-Video [96] Vicuna-7B [15] 860K 46.5 39.5 25.8 24.8 ShareGPT4Video [11] Llama3-8B [19] $\sim$1.0M 43.2 48.7 27.8 16.1 MiniCPM-V 2.6 [83] Qwen2-7B [80] $\sim$7.0M 57.6 66.8 40.0 45.2 LLaVA-OV (SI) [29] Qwen2-7B [80] $\sim$7.2M 60.3 61.8 31.4 36.4 LLaVA-OV [29] Qwen2-7B [80] $\sim$8.8M 48.0 71.7 33.2 45.6 LLaVA-Video [97] Qwen2-7B [80] $\sim$7.4M 45.1 67.0 34.1 53.2 InternVL2 [12] InternLM2.5-7B [7] N/A 58.6 66.5 39.8 45.5 Qwen2-VL-Instruct [74] Qwen2-7B [80] N/A 48.3 64.9 38.2 59.4 Qwen2-VL-Instruct [74] Qwen2-72B [80] N/A 55.5 74.7 45.5 74.6 Our models LLaVA-Next-Inst-IT Vicuna-7B [15] 920K 68.6 63.0 49.3 42.1 LLaVA-Next-Inst-IT Qwen2-7B [80] 920K 67.9 75.3 45.7 53.3
@@
Table 1: Performance of LMMs on Inst-IT Bench. We conduct extensive evaluations on Inst-IT Bench, including state-of-the-art open-source image models, video models, and cutting-edge proprietary models. #IT indicates the number of training samples used during the instruction-tuning stage. N/A indicates that the number of training samples is unknown.
@@ 4.1 Implementation Details
We use LLaVA-Next [41] as our baseline multi-modal large model due to its widespread adoption. In the default configuration, Vicuna-1.5-7B [15] serves as the language model with CLIP-ViT-336 [62] as the vision encoder. We utilize the AdamW [44] with a cosine learning rate schedule for optimization. During the vision-language alignment stage, we use the LCS-558K dataset [40], and for the supervised fine-tuning stage, we leverage the open-source LLaVA-Next-DATA [43]. For single images, we split the original image into up to 4 sub-images based on its aspect ratio following the AnyRes [41] approach, and then concatenate the global image with these sub-images. For multiple images and video inputs, we skip the AnyRes procedure and encode every single image. Additionally, we apply $2\times 2$ spatial pooling to reduce the number of visual tokens for video inputs. We limit the maximum number of frames to 32 and the context length of LLMs to 6K due to GPU and memory constraints. To enhance instance-level understanding with our Inst-IT Dataset, we combine Inst-IT Dataset with LLaVA-Next-DATA and introduce an additional continuous supervised fine-tuning (SFT) stage. In this stage, we freeze the first 12 layers of the vision encoder to mitigate potential distribution shifts caused by visually prompted images. Furthermore, we use Qwen2-7B [80] with SigLIP-SO400M-384 [89] for improved performance in our main experiment, and Qwen2-1.5B with CLIP-ViT-336 for efficiency in our ablation study.
Method LLM Vision Encoder AI2D [25] MMMU [87] POPE [34] GQA [24] MM-Vet [85]
@@
LLaVA-1.5 [40] Vicuna-7B [15] CLIP-ViT-Large [61] 54.8 35.3 85.9 62.0 30.5
LLaVA-Next [41] Vicuna-7B [15] CLIP-ViT-Large [61] 66.6 35.1 86.4 64.2 44.1
DeepStack-L [54] Vicuna-7B [15] CLIP-ViT-Large [61] - 35.7 86.7 63.1 29.9
DeepStack-L-HD [54] Vicuna-7B [15] CLIP-ViT-Large [61] - 35.6 86.5 65.2 37.5
VILA [38] Vicuna-7B [15] CLIP-ViT-Large [61] - - 85.5 62.3 34.9
ShareGPT4V [10] Vicuna-7B [15] CLIP-ViT-Large [61] - - - 37.6
MM-1.5 [91] MM-LLM-7B [91] MM-CLIP [91] 72.0 41.8 88.6 - 42.2
InternVL2 [12] InternLM-7B [68] InternViT-300M [12] 83.8 49.3 - - 60.0
LLaVA-OV (SI) [29] Qwen2-7B [80] SigLIP-SO400M [89] 81.6 47.3 - - 58.8
LLaVA-OV [29] Qwen2-7B [80] SigLIP-SO400M [89] 81.4 48.8 - - 57.5
Qwen2-VL-Instruct [74] Qwen2-7B [80] DFN-CLIP-H [20] 83.0 54.1 - - 62.0
LLaVA-Next-Inst-IT Vicuna-7B [15] CLIP-ViT-Large [61] 71.0 37.4 87.2 65.9 38.1
LLaVA-Next-Inst-IT Qwen2-7B [80] SigLIP-SO400 [89] 78.7 42.7 87.6 65.5 44.7
Table 2: Main results on image benchmarks.
Method LLM Vision Encoder ANet-QA [86] EgoSchema [47] Next-QA [77] VideoMME [21] TempCompass [42]
(open-ended) (subset) (val) (w/o subs) (3 avg)
DeepStack-L [54] Vicuna-7B [15] CLIP-ViT-Large [61] 49.3 38.4 61.0 - -
IG-VLM [26] Vicuna-7B [15] CLIP-ViT-Large [61] 54.3 35.8 63.1 - -
LLaVA-Next [41] Vicuna-7B [15] CLIP-ViT-Large [61] 53.8 - - - -
SF-LLaVA [78] Vicuna-7B [15] CLIP-ViT-Large [61] 55.5 47.2 64.2 -
Video-ChatGPT [46] Vicuna-7B [15] CLIP-ViT-Large [61] 35.2 47.3 - - -
VideoLLaMA2 [13] Vicuna-7B [15] CLIP-ViT-Large [61] 50.2 - 51.7 - -
LLaVA-Next-Video [96] Vicuna-7B [15] CLIP-ViT-Large [61] 53.5 43.9 - 46.5 -
LLaVA-Next-Video-DPO [96] Vicuna-7B [15] CLIP-ViT-Large [61] 60.2 - - 58.3
LongVA [93] Qwen2-7B [80] CLIP-ViT-Large [61] - 68.3 52.4 61.3
MM-1.5-Video-SFT [91] MM-LLM-7B [91] MM-CLIP [91] 60.9 57.2 76.8 53.5 -
InternVL2 [12] InternLM-7B [68] InternViT-300M [12] - - - 54.0 -
LLaVA-OV [29] Qwen2-7B [80] SigLIP-SO400M [89] 56.6 60.1 79.4 58.2 69.4
LLaVA-Video [97] Qwen2-7B [80] SigLIP-SO400M [89] 56.5 57.3 83.2 63.3 -
Qwen2-VL-Instruct [74] Qwen2-7B [80] DFN-CLIP-H [20] - 66.7 - 63.3 72.9
LLaVA-Next-Inst-IT Vicuna-7B [15] CLIP-ViT-Large [61] 53.7 57.8 70.2 44.3 59.8
LLaVA-Next-Inst-IT Qwen2-7B [80] SigLIP-SO400 [89] 55.2 50.4 73.0 54.0 63.9
Table 3: Main results on video benchmarks. We report the average of 3 parts (MCQA, Y/N, Caption Match) of TempCompass for determinism results.
@@ 4.2 Main Experiments
Results on Inst-IT Bench. We conduct extensive evaluate of existing multimodal models on Inst-IT Bench. The results in Tab.1, show that with instruction tuning on LLaVA-NeXT [41] using Inst-IT Dataset, the model achieves a significant improvement of nearly 20% on average score, validating the effectiveness of our Inst-IT Dataset. Moreover, although ViP-LLaVA [5] utilizes visual prompts for instruction tuning, it shows minor improvement over the baseline LLaVA-1.5 [40], possibly due to overfitting to its training data. In contrast, our model demonstrates consistent improvements on other instance understanding benchmarks, such as ViP-Bench [5], as well as on general-purpose evaluation sets like AI2D and Egoschema (further discussions are in the following sections). This suggests that model trained with Inst-IT Dataset can generalizes well to out-of-domain data. Besides, as Qwen2VL-72B does not show substantial improvements over its smaller 7B model, indicating that simply scaling up the model size cannot address the challenges in instance understanding. Similarly, by comparing the amount of instruction tuning data used by each models, we observe that large-scale coarse-grained annotations does not lead to essential improvements either. This highlights the importance of instance-specific annotated data, as we achieve impressive performance by training a 7B model on 920k data.
@@
Results on academic benchmarks. We evaluate our models using the LMMs-Eval [92] toolkit. To ensure a fair comparison with other models, we primarily report results from their original papers or from reproduced results in other studies. On widely-used image benchmarks shown in Tab. 2, our Inst-IT consistently outperforms LLaVA-Next, our direct baseline model. Furthermore, when utilizing a more advanced language model and a vision encoder, our method achieves performance comparable to large-scale SFT LMMs, such as LLaVA-OV and Qwen2-VL-Instruct, while requiring significantly less computational and data cost for instruction tuning. For video understanding benchmarks Tab. 3, our method significantly outperforms both LLaVA-Next and LLaVA-Next-Video. These consistent improvements demonstrate that enhancing instance-level understanding through explicit visual prompt instruction tuning is an effective strategy for improving generic spatiotemporal understanding.

### v2

@@ 3.3 Inst-IT Bench
Existing benchmarks primarily focus on global understanding, failing to provide more in-depth insights into the instance-level comprehension. We present the Inst-IT Bench, specifically designed to diagnose multimodal instance-level understanding in both images and videos.
@@
Construction process. To prevent data leakage, we use videos from the test split, ensuring no overlap with Inst-IT Dataset. We apply the pipeline in Sec.˜3.1 to generate 20 open-ended QA pairs for each image and video. Then, we manually review these QA pairs to ensure their accuracy, diversity, and difficulty. Overly simple questions are removed to ensure the remaining ones are instances-centric. We also refine the questions and answers, making necessary rephrasing to ensure correctness. After this rigorous checking process, each sample retains an average of 3.7 carefully polished QA pairs. In addition, we generate three hard negative for each question to construct a multiple-choice QA data with four options. More details are provided in Sec.˜B.1.
Statistics. Inst-IT Bench comprises 1,000 QA pairs for 338 images and 1,000 QA pairs for 206 videos. Each QA pair supports two evaluation formats, i.e. open-ended and multiple-choice.
@@
Metrics. For open-ended QAs, we leverage GPT-4o to evaluate the response from a model based on its similarity to the ground-truth answer. For multiple-choice QAs, we calculate the average accuracy across all samples. More details about the metric calculations can be found in Sec.˜B.2.
Comparison with existing instance understanding benchmarks. Tab.˜1 (right) highlights the main differences between Inst-IT Bench and existing instance understanding benchmarks such as RefCOCO [refcoco], RefCOCOg [refcocog] and ViP-Bench [vipllava]: (1) its inclusion of evaluation data for both images and videos, pioneered the evaluation in video LMMs; and (2) it supports both open-ended and multiple-choice formats, enabling comprehensive evaluation.
Table 1: Comparison of Inst-IT with existing datasets and instance understanding benchmarks. Left: Instruction tuning datasets. Right: Instance understanding benchmarks. Img and Vid indicate whether the data contains images or videos, respectively. Inst denotes the availability of instance-level annotations. OE and MC indicate open-ended and multiple-choice QA.
Img Vid Inst
ShareGPT4Video [sharegpt4video]
LLaVA-Video [llavavideo]
ViP-LLaVA-Data [vipllava]
[…]
Img Vid Task
RefCOCO [refcoco] caption
RefCOCOg [refcocog] caption
ViP-Bench [vipllava] OE
[…]
3.4 Instruction tuning with explicit visual prompt
@@
Architecture. We adopt the widely-used LLaVA-NeXT [llavanext] architecture to evaluate the effectiveness of our Inst-IT. We train our model under an image-video joint training pipeline, where we mix our Inst-IT Dataset with the open-source LLaVA-NeXT-DATA [llavanext-data]. For single-image samples, we follow the original AnyRes paradigm [llavanext] to split and encode sub-images according to the aspect ratio. For video and multi-image data, we batch the samples together, encode them, and flatten them into a sequence. Additionally, we apply $2\times 2$ spatial pooling to reduce the number of visual tokens in the video inputs. More details are in Sec.˜4.1.
Converting Inst-IT Dataset into instruction tuning format. Inst-IT Dataset provides annotations at multiple levels of granularity. For the instance- and image-level captions in figure˜2(a), we use a single frame as input and structure the task as a two-turn dialogue: the model is first prompted to describe all individual instances, followed by a holistic description of the entire scene. To capture temporal dynamics, we use temporal captions from figure˜2(a), asking the model to describe the differences between two consecutive frames. The video-level description in figure˜2(b) is treated as a captioning task, where the model is instructed to generate a summary based on all video frames. For the open-ended QA pairs in figure˜2(c), we organize them into a multi-turn conversation, with the model answering one question per turn. In total, we construct 243k instruction tuning samples in the form of single-turn and multi-turn dialogues. All images and video frames are augmented with SoM visual prompts to explicitly provide instance-level guidance.
@@
Table 2: Results on Inst-IT Bench. We conduct evaluations on Inst-IT Bench, including state-of-the-art open-source image models, video models, and cutting-edge proprietary models. #IT indicates the number of training samples used during the instruction-tuning stage. N/A indicates that the number is unknown. OE and MC represent open-ended and multiple-choice evaluations, respectively.
Model LLM #IT Image Video
OE Q&A MC Q&A OE Q&A MC Q&A
Random Guess - N/A - 25.0 - 25.0
GPT-4o [openai2024gpt4o] - N/A 74.1 84.8 65.5 81.0
Gemini-1.5-flash [Reid2024Gemini1U] - N/A 65.3 79.5 57.9 75.8
Open-source image models
LLaVA-1.5 [llava1d5] Vicuna-7B 665K 41.6 32.1 - -
ViP-LLaVA [vipllava] Vicuna-7B $\sim$1.2M 42.1 29.2 - -
SoM-LLaVA [somllava] Vicuna-7B 695K 45.1 40.0 - -
[…]
Open-source video models
LLaVA-NeXT-Video [llavanext-video] Vicuna-7B 860K 46.5 39.5 25.8 24.8
ShareGPT4Video [sharegpt4video] Llama3-8B $\sim$1.0M 43.2 48.7 27.8 16.1
LLaVA-OV (SI) [li2024llava] Qwen2-7B $\sim$7.2M 60.3 61.8 31.4 36.4
LLaVA-OV [li2024llava] Qwen2-7B $\sim$8.8M 48.0 71.7 33.2 45.6
LLaVA-Video [llavavideo] Qwen2-7B $\sim$7.4M 45.1 67.0 34.1 53.2
InternVL2 [internvl] InternLM2.5-7B N/A 58.6 66.5 39.8 45.5
Qwen2-VL-Instruct [Qwen2VL] Qwen2-72B N/A 55.5 74.7 45.5 74.6
Our models
[…]
@@ 4.1 Implementation details
[…]
Table 3: Main results on image benchmarks.
Method LLM Vision Encoder AI2D[kembhavi2016diagram] MMMU[yue2024mmmu] POPE[li2023evaluating] GQA[hudson2019gqa] ChartQA[masry2022chartqa]
@@
LLaVA-1.5 [llava1d5] Vicuna-7B CLIP-ViT-Large 54.8 35.3 85.9 62.0 18.2
DeepStack-L [deepstack] Vicuna-7B CLIP-ViT-Large - 35.7 86.7 63.1 21.0
DeepStack-L-HD [deepstack] Vicuna-7B CLIP-ViT-Large - 35.6 86.5 65.2 56.3
VILA [lin2024vila] Vicuna-7B CLIP-ViT-Large - - 85.5 62.3 -
LLaVA-OV (SI) [li2024llava] Qwen2-7B SigLIP-SO400M 81.6 47.3 - - 78.8
LLaVA-OV [li2024llava] Qwen2-7B SigLIP-SO400M 81.4 48.8 - - 80.0
Qwen2-VL-Instruct [Qwen2VL] Qwen2-7B DFN-CLIP-H 83.0 54.1 - - 83.0
[…]
Table 4: Main results on video benchmarks. We report the average of MCQA, Y/N and CM in TempCompass for determinism results. indicates results reproduced by us.
Method LLM Vision Encoder ANetQA[yu2019activityqa] EgoSchema[mangalam2023egoschema] NExTQA[xiao2021next] VideoMME[fu2024video] TempCompass[liu2024tempcompass]
(oe) (subset) (mc) (w/o subs) (3 avg)
DeepStack-L [deepstack] Vicuna-7B CLIP-ViT-Large 49.3 38.4 61.0 - -
Video-ChatGPT [video-chatgpt] Vicuna-7B CLIP-ViT-Large 35.2 47.3 - - -
VideoLLaMA2 [cheng2024videollama] Vicuna-7B CLIP-ViT-Large 50.2 - 51.7 - -
LLaVA-Next-Video [llavanext-video] Vicuna-7B CLIP-ViT-Large 53.5 43.9 - 46.5 -
InternVL2 [internvl] InternLM-7B InternViT-300M - - - 54.0 -
LLaVA-OV [li2024llava] Qwen2-7B SigLIP-SO400M 56.6 60.1 79.4 58.2 69.4
LLaVA-Video [llavavideo] Qwen2-7B SigLIP-SO400M 56.5 57.3 83.2 63.3 -
Qwen2-VL-Instruct [Qwen2VL] Qwen2-7B DFN-CLIP-H - 66.7 - 63.3 72.9
[…]
@@ 4.2 Main experiments
Results on Inst-IT Bench. We conduct extensive evaluations on Inst-IT Bench. The results in Tab.˜2 show that with instruction tuning using Inst-IT Dataset, our models achieve a significant improvement of nearly 20% on average score, validating the effectiveness of Inst-IT. Moreover, although ViP-LLaVA [vipllava] utilizes visual prompts for instruction tuning, it shows minor improvement over its baseline, i.e. LLaVA-1.5 [llava1d5], possibly due to overfitting to its training data. In contrast, our model demonstrates consistent improvements on other instance understanding benchmarks, such as ViP-Bench [vipllava] and RefCOCOg [refcocog] (Sec.˜4.3), as well as on general-purpose evaluation sets like AI2D and Egoschema (will be discussed in the following sections). This suggests that the model trained with Inst-IT generalizes well to other tasks. Qwen2VL-72B does not show substantial improvements over its smaller 7B model, indicating that simply scaling up the model size cannot address the challenges in instance understanding. Similarly, by comparing the amount of instruction tuning data used by each model, we observe that large-scale coarse-grained annotations do not lead to essential improvements either. This highlights the importance of instance-specific annotated data.
@@
Results on generic benchmarks. To evaluate general understanding capabilities, we assess our models on several widely used image and video benchmarks using the LMMs-Eval [zhang2024lmms]. To ensure a fair comparison with other models, we primarily report results from their original papers or reproduced results in previous studies. On generic image benchmarks, as shown in Tab.˜3, Inst-IT consistently outperforms our direct baseline model, i.e. LLaVA-NeXT. The improvement in AI2D, a benchmark that requires grounding and referring understanding capability, is particularly clear. This suggests that Inst-IT effectively boosts the model in fine-grained understanding. Furthermore, when utilizing a more advanced language model and vision encoder, our method achieves performance comparable to large-scale SFT LMMs, such as LLaVA-OV and Qwen2-VL-Instruct, while requiring significantly less computational and data cost. For video understanding benchmarks in Tab.˜4, Inst-IT significantly outperforms both LLaVA-NeXT and LLaVA-NeXT-Video. These consistent improvements demonstrate that enhancing instance-level understanding through explicit visual prompted instruction tuning is an effective strategy for improving generic spatiotemporal understanding capabilities.
Table 5: Results on ViP-Bench. We perform evaluation with our Inst-IT models without fine-tuning.
Model Synthesized visual prompts Visual prompts from human
Rec OCR Know Math Rel Lang All Rec OCR Know Math Rel Lang All
[…]
InstructBLIP-7B [instructblip] 36.9 16.3 34.2 22.3 26.8 7.5 31.7 38.9 17 35.4 9.7 29.3 17.5 33.3
Shikra-7B [chen2023shikra] 40.2 10.0 28.0 3.5 18.9 20.6 33.7
GPT4ROI-7B [zhang2023gpt4roi] 35.6 16.7 29.7 9.7 32.5 13.8 35.1
Kosmos-2 [peng2023kosmos] 29.5 14.2 18.5 9.7 7.5 21.9 26.9
LLaVA-1.5-7B [llava1d5] 50.8 12.4 49.2 6.5 51.8 23.8 41.6 49.1 13.0 42.9 9.7 50.0 27.5 40.2
Qwen-VL-Chat [bai2023qwen] 43.0 30.4 40.2 9.7 25.7 28.7 39.2 48.7 22.1 41.2 6.5 48.2 25.0 41.7
ViP-LLaVA-7B [vipllava] 54.8 18.8 52.9 9.7 53.9 42.5 45.5 55.3 17.6 45.9 8.1 44.6 33.1 46.8
[…]
4.3 Evaluation on other instance-understanding benchmarks
To assess whether our model has learned generalizable instance understanding capability, we conducted evaluations on out-of-domain instance understanding benchmarks in zero-shot manner.